# On the Generalization of Handwritten Text Recognition Models

### v1

@@ Abstract
Recent advances in Handwritten Text Recognition (HTR) have led to significant reductions in transcription errors on standard benchmarks under the i.i.d. assumption, thus focusing on minimizing in-distribution (ID) errors. However, this assumption does not hold in real-world applications, which has motivated HTR research to explore Transfer Learning and Domain Adaptation techniques. In this work, we investigate the unaddressed limitations of HTR models in generalizing to out-of-distribution (OOD) data. We adopt the challenging setting of Domain Generalization, where models are expected to generalize to OOD data without any prior access. To this end, we analyze 336 OOD cases from eight state-of-the-art HTR models across seven widely used datasets, spanning five languages. Additionally, we study how HTR models leverage synthetic data to generalize. We reveal that the most significant factor for generalization lies in the textual divergence between domains, followed by visual divergence. We demonstrate that the error of HTR models in OOD scenarios can be reliably estimated, with discrepancies falling below 10 points in 70% of cases. We identify the underlying limitations of HTR models, laying the foundation for future research to address this challenge.
I Introduction
@@
Figure 1: Average performance of HTR models for in-distribution (ID) and out-of-distribution (OOD) scenarios. The significant performance drop in OOD scenarios highlights the limited generalization capability of current models. Details in Sect. IV.
@@
II Related Work
II-A Handwritten Text Recognition
@@ III-B Baseline HTR Models
CTC. Most architectures in the modern HTR literature employ the CTC objective [49] as an effective method for aligning images with unsegmented transcriptions. We selected four distinct architectures with different feature extractors. These architectures are: CRNN [10], VAN [11], CNN-SAN [78], HTR-VIT [15]. This selection comprises different CTC-based architectures with a broad range of encoders and decoders.
Sequence-to-sequence. Sequence-to-sequence models (Seq2Seq) have been dominating the arena of language processing [50], particularly boosted by the introduction of the Transformer [55].111 In this work, we do not include TrOCR [16], as it was not originally designed for HTR but rather as a secondary application. To investigate this type of Seq2Seq architectures, we utilize the Transformer presented in [79] (ResNet + Transformer) since other pure Transformer-based architectures merely differ in post-processing techniques [64].
Hybrid. Several studies have adopted hybrid approaches for HTR [54, 13, 14]. Inspired by [80], which combined CTC for the encoder and Cross-Entropy (CE) for the decoder as a multitask learning for Speech Recognition, a similar method was first applied in HTR in [54]. This approach utilizes a hybrid loss function that merges CTC and CE losses as $\mathcal{L}=\lambda\mathcal{L_{\text{ctc}}}+(1-\lambda)\mathcal{L_{\text{ce}}},\lambda\in[0,1]$,
@@
Table I summarizes the architectures considered in this work.
TABLE I: Description of HTR architectures considered, including alignment type, number of parameters, and input image sizes. C=CNN; FCN=Fully Convolutional Network; SA=Self Attention; T=Transformer; Att=“Classic” attention.
@@
III-C Experimental setup
Data. We assessed the performance of the models in the following datasets: IAM [1], Rimes [81], Bentham [82], Saint-Gall (S.G.) [83], George Washington (G.W.) [3], Rodrigo [84] and ICFHR 2016 (READ 2016) [85]. Table II details each dataset.
TABLE II: Handwritten document datasets considered, with an indication of the language, historical period, number of writers, number of samples in each partition, and size of the alphabet ($\Sigma$).
@@ G.W. En 1755 1 325 168 163 68
Performance metrics. As the majority of works in the field, we assessed the performance of the models using the common Character Error Rate (CER). [^2] Results with the Word Error Rate will also be reported in Appendix VII. This ensures consistency with existing research and facilitates a more direct comparison with our findings. Additionally, we computed the Expected Calibration Error (ECE) for each model, following the definition in [86]. We aim to assess not only the accuracy of the models but also their capacity to generate calibrated probability outputs. This is critical for understanding model behavior and uncertainty when applied to unseen OOD data, thereby offering a more comprehensive evaluation of model performance.
Implementation details. Each model was trained from scratch for 500 epochs, with the best model saved based on validation performance (CER) for the corresponding domain. Training is always performed with a single source domain, and the run is stopped if the CER on the validation set does not improve at least 0.1 within 100 epochs. We implemented a comprehensive set of typical data augmentation techniques from HTR literature [87, 88]: rotation, dilation, erosion, random perspective, elastic transformation, shearing, and Gaussian noise, applied with a probability of 0.5. Lastly, we grayscaled all images as in [10]. Hyperparameters such as architecture, input size, batch size, optimizers, and schedulers were consistent with the original configurations. Further details on data augmentation and hyperparameters are provided in Appendix VIII. Given that each dataset contains a unique alphabet, we combined all alphabets using Unicode [^3] https://pypi.org/project/Unidecode/ to standardize evaluations, resulting in a vocabulary of 94 characters including special symbols for the beginning of sequence [BOS], padding [PAD], unknown [UNK], and the end of sequence [EOS]. No post-processing techniques or lexicon-based predictions were applied to the outputs of the models.
TABLE III: In-distribution (ID) and out-of-distribution (OOD) results (CER %) for HTR models across datasets. The OOD result is reported from the best-performing source. Results marked with $\ast$ indicate outliers, meaning that the model did not converge in the ID setting. Average results (bottom row) are computed filtering out outliers. $\dagger$ denotes architectures implemented from the papers (no code provided).
@@
IAM 6.4     34.9 6.6     28.6 15.0     31.5 5.8     33.7 8.0     42.1 7.5     49.1 7.9     42.0 8.9     41.3
Rimes 3.7     25.0 5.6     21.3 12.0     29.8 7.9     28.3 5.7     32.0 6.9     35.5 5.0     30.8 5.1     29.4
G.W. 8.2     31.1 9.3     32.0 9.0     49.8 34.9    38.6 78.4    44.0 53.8    43.6 79.6    32.3 25.2    32.1
Bentham 4.7     25.3 7.4     26.6 10.0     38.9 8.4     33.3 8.5     39.4 8.5     43.5 6.0     33.8 6.1     33.3
S.G. 7.2     33.6 7.8     39.8 8.6     35.0 17.1     36.5 78.7    51.8 76.9    55.3 12.5     37.8 9.2     38.7
Rodrigo 1.7     40.9 2.3     38.5 3.8     45.2 3.9     38.5 2.6     60.6 3.8     65.3 2.0     48.4 2.2     47.4
$\text{ICFHR}_{\text{2016}}$ 5.2     78.7 7.5     75.3 17.0     83.4 11.6     79.6 7.8     92.6 9.5     85.2 5.9     90.5 6.0     85.1
Average 5.3     38.5 6.7     37.4 9.7     44.8 7.5     41.2 6.5     51.8 7.2     53.9 6.5     45.1 6.3     43.9
IV Practical out-of-distribution insights
@@
IV-A HTR performance
The results of the first experiments are reported in Table III, where the rows represent the target domain, the first column indicates the ID error, and the OOD column shows the best error achieved from any other source domain. In this scenario, the typical ID model selection is applied by choosing the one with the best validation error in the source domain. The average performance of both models in the two scenarios (ID, OOD) is presented in the last row, with the best results highlighted in bold.
OOD results are terrible. OOD results are poor across all architectures and alignment types, with CER values ranging from 37.4% to 53.9%—error rates that are notably high for a transcription system. The model that performs best in terms of generalization is VAN, showing only a slight absolute improvement of 1% over the second-best model (CRNN), which also uses CTC alignment. When comparing the best model (VAN) to the worst model (Michael) in terms of generalization, the absolute improvement is only 16.5%. The average difference between ID and OOD CERs is 37.6%, posing that generalization gaps in HTR remain significant even under the best OOD conditions.
Alignment type does not really matter, but choose CTC. OOD performance is ordered and color-coded according to the alignment type in Fig. 3. While the observed differences are not substantial, a macro-level analysis reveals that models utilizing CTC alignment exhibit the best generalization performance, followed by hybrid approaches, and finally, purely Seq2Seq models. The results suggest that, within the DG scenario, a CTC-alignment model is the most effective choice. Notably, even the model with the poorest OOD performance (C-SAN) is comparable to the second-best alignment type (hybrid approach, VLT), with CER performances of 44.8% and 43.9%, respectively. Grouping models by their decoding method—CTC or autoregressive (AR)—we observe that CTC-based models achieve an average CER of 40.4, while AR models have a CER of 48.7, resulting in a relative drop of 20%.
@@ IV-A HTR performance
IV-B Leveraging synthetic data
This section investigates how HTR models utilize synthetic data for predicting on the typical HTR datasets. The study has two primary objectives: (1) to assess the effectiveness of each architecture in leveraging synthetic data relative to real data, and (2) to quantify the influence of the underlying language model (LM) on the overall performance. We designed controlled experiments to evaluate the effect of synthetic text generation across multiple languages on real datasets’ performances. We generated synthetic 100K text lines randomly selected from WIT dataset [89] in English, French, Spanish, German, and Latin using 4,000 publicly available fonts [^4] Fonts were sourced from open-access repositories like 1001fonts.com. with handwritten styles ensuring an equal number of lines for each language. We excluded HTG methods [20, 21, 18] as they rely on labeled data targeting ID performance, limiting their applicability to DG. For this scenario, we compare the performance of models trained on synthetic datasets rendered in the same language as the target versus those trained on a different language. To ensure fair comparisons, we consistently selected the best-performing models based on validation results for the target dataset regarding the language used. The results are shown in Fig. 4, where results are sorted from best to worst (left-to-right) to facilitate visualization.
@@ IV-B Leveraging synthetic data
Synthetic data (slightly) improves OOD, even blindly. The first observation in Fig. 4 is that using synthetic data consistently improves OOD performance compared to real data, even when a different language is used from the target domain. Although the CER remains high, ranging from 32.9 to 45.7, models generalize an average of 6 points better when using synthetic data. In the best case, where the target domain language is known, models show an absolute gain of over 11 CER points.
Choose CTC again (AR-models are more biased). Fig. 4 shows that the model that best utilizes synthetic data is the CRNN architecture, followed by VAN. While the differences are less pronounced compared to the OOD scenario using real data, the conclusion remains the same: CTC alignment is still the best option. When grouping models by their decoding method, whether CTC or autoregressive (AR), we find that the relative decrease compared to training with synthetic data—same or a different language—is 25% for AR and 19% for CTC. This intuitively reflects the nature and bias of AR models: they are explicitly trained to compute an LM, unlike CTC models, making the latter more robust to language variations.
IV-C Model selection and capacity
In this section, we evaluate the performance of HTR models through two key factors from the DG literature: model selection [40] and model complexity [41]. We compare three OOD model selection strategies inspired by DG studies: (1) No-selection (ID selection), where the best model is chosen based on source validation performance (standard i.i.d. setting, as shown in Table III) used as a baseline; (2) Heldout selection, where model selection is based on the average performance across all validation data sources excluding the target dataset [40]; (3) Oracle-based selection, which sets an upper performance bound by selecting the best model based on validation results from the target dataset [40]. As with prior analyses, OOD performance reflects the best result, regardless of the source domain used for training. We present the results in Fig. 5, where the capacity of the HTR models (measured in millions of parameters) is compared to their OOD performance across the three model selection methods using real and synthetic data.

### v2

@@ Abstract
Recent advances in Handwritten Text Recognition (HTR) have led to significant reductions in transcription errors on standard benchmarks under the i.i.d. assumption, thus focusing on minimizing in-distribution (ID) errors. However, this assumption does not hold in real-world applications, which has motivated HTR research to explore Transfer Learning and Domain Adaptation techniques. In this work, we investigate the unaddressed limitations of HTR models in generalizing to out-of-distribution (OOD) data. We adopt the challenging setting of Domain Generalization, where models are expected to generalize to OOD data without any prior access. To this end, we analyze 336 OOD cases from eight state-of-the-art HTR models across seven widely used datasets, spanning five languages. Additionally, we study how HTR models leverage synthetic data to generalize. We reveal that the most significant factor for generalization lies in the textual divergence between domains, followed by visual divergence. We demonstrate that the error of HTR models in OOD scenarios can be reliably estimated, with discrepancies falling below 10 points in 70% of cases. We identify the underlying limitations of HTR models, laying the foundation for future research to address this challenge. Code is available at github.com/carlos10garrido/HTR-OOD.
1 Introduction
Figure 1: Average performance of HTR models for in-distribution (ID) and out-of-distribution (OOD) scenarios. The significant performance drop in OOD scenarios highlights the limited generalization capability of current models. Details in Sect. 4.
@@ 1 Introduction
2 Related Work
2.1 Handwritten Text Recognition
@@ 3.2 Baseline HTR Models
CTC.
Most architectures in the modern HTR literature employ the CTC objective [32] as an effective method for aligning images with unsegmented transcriptions. We selected four distinct architectures with different feature extractors. These architectures are: CRNN [64], VAN [20], CNN-SAN [26], HTR-VIT [47]. This selection comprises different CTC-based architectures with a broad range of encoders and decoders.
Sequence-to-sequence.
Sequence-to-sequence models (Seq2Seq) have been dominating the arena of language processing [9], particularly boosted by the introduction of the Transformer [83].111 In this work, we do not include TrOCR [46], as it was not originally designed for HTR but rather as a secondary application. To investigate this type of Seq2Seq architectures, we utilize the Transformer presented in [37] (ResNet + Transformer) since other pure Transformer-based architectures merely differ in post-processing techniques [90].
Hybrid.
Several studies have adopted hybrid approaches for HTR [54, 10, 11]. Inspired by [42], which combined CTC for the encoder and Cross-Entropy (CE) for the decoder as a multitask learning for Speech Recognition, a similar method was first applied in HTR in [54]. This approach utilizes a hybrid loss function that merges CTC and CE losses as $\mathcal{L}=\lambda\mathcal{L_{\text{ctc}}}+(1-\lambda)\mathcal{L_{\text{ce}}},\lambda\in[0,1]$,
@@
Table 1 summarizes the architectures considered in this work.
Table 1: Description of HTR architectures considered, including alignment type, number of parameters, and input image sizes. C=CNN; FCN=Fully Convolutional Network; SA=Self Attention; T=Transformer; Att=“Classic” attention.
@@
3.3 Experimental setup
Data.
We assessed the performance of the models in the following datasets: IAM [52], Rimes [41], Bentham [19], Saint-Gall (S.G.) [71], George Washington (G.W.) [66], Rodrigo [72] and ICFHR 2016 (READ 2016) [77]. Table 2 details each dataset.
Table 2: Handwritten document datasets considered, with an indication of the language, historical period, number of writers, number of samples in each partition, and size of the alphabet ($\Sigma$).
@@ G.W. En 1755 1 325 168 163 68
Performance metrics.
As the majority of works in the field, we assessed the performance of the models using the common Character Error Rate (CER). [^2] Results with the Word Error Rate will also be reported in Appendix 7. This ensures consistency with existing research and facilitates a more direct comparison with our findings. Additionally, we computed the Expected Calibration Error (ECE) for each model, following the definition in [7]. We aim to assess not only the accuracy of the models but also their capacity to generate calibrated probability outputs. This is critical for understanding model behavior and uncertainty when applied to unseen OOD data, thereby offering a more comprehensive evaluation of model performance.
Implementation details.
Each model was trained from scratch for 500 epochs, with the best model saved based on validation performance (CER) for the corresponding domain. Training is always performed with a single source domain, and the run is stopped if the CER on the validation set does not improve at least 0.1 within 100 epochs. We implemented a comprehensive set of typical data augmentation techniques from HTR literature [51, 68]: rotation, dilation, erosion, random perspective, elastic transformation, shearing, and Gaussian noise, applied with a probability of 0.5. Lastly, we grayscaled all images as in [64]. Hyperparameters such as architecture, input size, batch size, optimizers, and schedulers were consistent with the original configurations. Further details on data augmentation and hyperparameters are provided in Appendix 8. Given that each dataset contains a unique alphabet, we combined all alphabets using Unicode [^3] https://pypi.org/project/Unidecode/ to standardize evaluations, resulting in a vocabulary of 94 characters including special symbols for the beginning of sequence [BOS], padding [PAD], unknown [UNK], and the end of sequence [EOS]. No post-processing techniques or lexicon-based predictions were applied to the outputs of the models.
Table 3: In-distribution (ID) and out-of-distribution (OOD) results (CER %) for HTR models across datasets. The OOD result is reported from the best-performing source. Results marked with $\ast$ indicate outliers, meaning that the model did not converge in the ID setting. Average results (bottom row) are computed filtering out outliers. $\dagger$ denotes architectures implemented from the papers (no code provided).
@@
IAM 6.4 34.9${}^{\text{{\color[rgb]{1,0,0}(+28.5)}}}$ 6.6 28.6${}^{\text{{\color[rgb]{1,0,0}(+22.0)}}}$ 15.0 31.5${}^{\text{{\color[rgb]{1,0,0}(+16.6)}}}$ 5.8 33.7${}^{\text{{\color[rgb]{1,0,0}(+27.9)}}}$ 8.0 42.1${}^{\text{{\color[rgb]{1,0,0}(+34.1)}}}$ 7.5 49.1${}^{\text{{\color[rgb]{1,0,0}(+41.6)}}}$ 7.9 42.0${}^{\text{{\color[rgb]{1,0,0}(+34.1)}}}$ 8.9 41.3${}^{\text{{\color[rgb]{1,0,0}(+32.4)}}}$
Rimes 3.7 25.0${}^{\text{{\color[rgb]{1,0,0}(+21.2)}}}$ 5.6 21.3${}^{\text{{\color[rgb]{1,0,0}(+15.6)}}}$ 12.0 29.8${}^{\text{{\color[rgb]{1,0,0}(+19.8)}}}$ 7.9 28.3${}^{\text{{\color[rgb]{1,0,0}(+20.4)}}}$ 5.7 32.0${}^{\text{{\color[rgb]{1,0,0}(+26.3)}}}$ 6.9 35.5${}^{\text{{\color[rgb]{1,0,0}(+28.6)}}}$ 5.0 30.8${}^{\text{{\color[rgb]{1,0,0}(+25.8)}}}$ 5.1 29.4${}^{\text{{\color[rgb]{1,0,0}(+24.3)}}}$
G.W. 8.2 31.1${}^{\text{{\color[rgb]{1,0,0}(+22.9)}}}$ 9.3 32.0${}^{\text{{\color[rgb]{1,0,0}(+22.7)}}}$ 9.0 49.8${}^{\text{{\color[rgb]{1,0,0}(+40.8)}}}$ 34.9 ${}^{\text{{\color[rgb]{1,0,0}}}}$ 78.4 ${}^{\text{{\color[rgb]{1,0,0}}}}$ 53.8 ${}^{\text{{\color[rgb]{1,0,0}}}}$ 79.6 ${}^{\text{{\color[rgb]{1,0,0}}}}$ 25.2 32.1${}^{\text{{\color[rgb]{1,0,0}(+6.9)}}}$
Bentham 4.7 25.3${}^{\text{{\color[rgb]{1,0,0}(+20.6)}}}$ 7.4 26.6${}^{\text{{\color[rgb]{1,0,0}(+19.2)}}}$ 10.0 38.9${}^{\text{{\color[rgb]{1,0,0}(+28.9)}}}$ 8.4 33.3${}^{\text{{\color[rgb]{1,0,0}(+24.9)}}}$ 8.5 39.4${}^{\text{{\color[rgb]{1,0,0}(+30.9)}}}$ 8.5 43.5${}^{\text{{\color[rgb]{1,0,0}(+34.9)}}}$ 6.0 33.8${}^{\text{{\color[rgb]{1,0,0}(+27.8)}}}$ 6.1 33.3${}^{\text{{\color[rgb]{1,0,0}(+27.2)}}}$
S.G. 7.2 33.6${}^{\text{{\color[rgb]{1,0,0}(+26.3)}}}$ 7.8 39.8${}^{\text{{\color[rgb]{1,0,0}(+32.0)}}}$ 8.6 35.0${}^{\text{{\color[rgb]{1,0,0}(+26.4)}}}$ 17.1 36.5${}^{\text{{\color[rgb]{1,0,0}(+19.3)}}}$ 78.7 ${}^{\text{{\color[rgb]{1,0,0}}}}$ 76.9 ${}^{\text{{\color[rgb]{1,0,0}}}}$ 12.5 37.8${}^{\text{{\color[rgb]{1,0,0}(+25.3)}}}$ 9.2 38.7${}^{\text{{\color[rgb]{1,0,0}(+29.4)}}}$
Rodrigo 1.7 40.9${}^{\text{{\color[rgb]{1,0,0}(+39.3)}}}$ 2.3 38.5${}^{\text{{\color[rgb]{1,0,0}(+36.2)}}}$ 3.8 45.2${}^{\text{{\color[rgb]{1,0,0}(+41.4)}}}$ 3.9 38.5${}^{\text{{\color[rgb]{1,0,0}(+34.6)}}}$ 2.6 60.6${}^{\text{{\color[rgb]{1,0,0}(+57.9)}}}$ 3.8 65.3${}^{\text{{\color[rgb]{1,0,0}(+61.5)}}}$ 2.0 48.4${}^{\text{{\color[rgb]{1,0,0}(+46.4)}}}$ 2.2 47.4${}^{\text{{\color[rgb]{1,0,0}(+45.3)}}}$
$\text{ICFHR}_{\text{2016}}$ 5.2 78.7${}^{\text{{\color[rgb]{1,0,0}(+73.5)}}}$ 7.5 75.3${}^{\text{{\color[rgb]{1,0,0}(+67.8)}}}$ 17.0 83.4${}^{\text{{\color[rgb]{1,0,0}(+66.4)}}}$ 11.6 79.6${}^{\text{{\color[rgb]{1,0,0}(+68.0)}}}$ 7.8 92.6${}^{\text{{\color[rgb]{1,0,0}(+84.8)}}}$ 9.5 85.2${}^{\text{{\color[rgb]{1,0,0}(+75.7)}}}$ 5.9 90.5${}^{\text{{\color[rgb]{1,0,0}(+84.6)}}}$ 6.0 85.1${}^{\text{{\color[rgb]{1,0,0}(+79.1)}}}$
Average 5.3 38.5${}^{\text{{\color[rgb]{1,0,0}(+33.2)}}}$ 6.7 37.4${}^{\text{{\color[rgb]{1,0,0}(+30.8)}}}$ 9.7 44.8${}^{\text{{\color[rgb]{1,0,0}(+35.4)}}}$ 7.5 41.2${}^{\text{{\color[rgb]{1,0,0}(+33.7)}}}$ 6.5 51.8${}^{\text{{\color[rgb]{1,0,0}(+45.3)}}}$ 7.2 53.9${}^{\text{{\color[rgb]{1,0,0}(+46.7)}}}$ 6.5 45.1${}^{\text{{\color[rgb]{1,0,0}(+38.5)}}}$ 6.3 43.9${}^{\text{{\color[rgb]{1,0,0}(+37.6)}}}$
4 Practical out-of-distribution insights
@@
4.1 HTR performance
The results of the first experiments are reported in Table 3, where the rows represent the target domain, the first column indicates the ID error, and the OOD column shows the best error achieved from any other source domain. In this scenario, the typical ID model selection is applied by choosing the one with the best validation error in the source domain. The average performance of both models in the two scenarios (ID, OOD) is presented in the last row, with the best results highlighted in bold. We compare three state-of-the-art VLMs [48, 59, 22] and TrOCR [46] in a zero-shot setting in Appendix 7, highlighting their limitations for HTR.
OOD results are terrible.
OOD results are poor across all architectures and alignment types, with CER values ranging from 37.4% to 53.9%—error rates that are notably high for a transcription system. The model that performs best in terms of generalization is VAN, showing only a slight absolute improvement of 1% over the second-best model (CRNN), which also uses CTC alignment. When comparing the best model (VAN) to the worst model (Michael) in terms of generalization, the absolute improvement is only 16.5%. The average difference between ID and OOD CERs is 37.6%, posing that generalization gaps in HTR remain significant even under the best OOD conditions.
Alignment type does not really matter, but choose CTC.
OOD performance is ordered and color-coded according to the alignment type in Fig. 3. While the observed differences are not substantial, a macro-level analysis reveals that models utilizing CTC alignment exhibit the best generalization performance, followed by hybrid approaches, and finally, purely Seq2Seq models. The results suggest that, within the DG scenario, a CTC-alignment model is the most effective choice. Notably, even the model with the poorest OOD performance (C-SAN) is comparable to the second-best alignment type (hybrid approach, VLT), with CER performances of 44.8% and 43.9%, respectively. Grouping models by their decoding method—CTC or autoregressive (AR)—we observe that CTC-based models achieve an average CER of 40.4, while AR models have a CER of 48.7, resulting in a relative drop of 20%.
@@ 4.1 HTR performance
4.2 Leveraging synthetic data
This section investigates how HTR models utilize synthetic data for predicting on the typical HTR datasets. The study has two primary objectives: (1) to assess the effectiveness of each architecture in leveraging synthetic data relative to real data, and (2) to quantify the influence of the underlying language model (LM) on the overall performance. We designed controlled experiments to evaluate the effect of synthetic text generation across multiple languages on real datasets’ performances. We generated synthetic 100K text lines randomly selected from WIT dataset [75] in English, French, Spanish, German, and Latin using 4,000 publicly available fonts [^4] Fonts were sourced from open-access repositories like 1001fonts.com. with handwritten styles ensuring an equal number of lines for each language. We excluded HTG methods [82, 17, 61] as they rely on labeled data targeting ID performance, limiting their applicability to DG. For this scenario, we compare the performance of models trained on synthetic datasets rendered in the same language as the target versus those trained on a different language. To ensure fair comparisons, we consistently selected the best-performing models based on validation results for the target dataset regarding the language used. The results are shown in Fig. 4, where results are sorted from best to worst (left-to-right) to facilitate visualization. We also report the full results with synthetic data in Appendix 7 (Table 7).
@@ 4.2 Leveraging synthetic data
Synthetic data (slightly) improves OOD, even blindly.
The first observation in Fig. 4 is that using synthetic data consistently improves OOD performance compared to real data, even when a different language is used from the target domain. Although the CER remains high, ranging from 32.9 to 45.7, models generalize an average of 6 points better when using synthetic data. In the best case, where the target domain language is known, models show an absolute gain of over 11 CER points.
Choose CTC again (AR-models are more biased).
Fig. 4 shows that the model that best utilizes synthetic data is the CRNN architecture, followed by VAN. While the differences are less pronounced compared to the OOD scenario using real data, the conclusion remains the same: CTC alignment is still the best option. When grouping models by their decoding method, whether CTC or autoregressive (AR), we find that the relative decrease compared to training with synthetic data—same or a different language—is 25% for AR and 19% for CTC. This intuitively reflects the nature and bias of AR models: they are explicitly trained to compute an LM, unlike CTC models, making the latter more robust to language variations.
4.3 Model selection and capacity
In this section, we evaluate the performance of HTR models through two key factors from the DG literature: model selection [33] and model complexity [29]. We compare three OOD model selection strategies inspired by DG studies: (1) No-selection (ID selection), where the best model is chosen based on source validation performance (standard i.i.d. setting, as shown in Table 3) used as a baseline; (2) Heldout selection, where model selection is based on the average performance across all validation data sources excluding the target dataset [33]; (3) Oracle-based selection, which sets an upper performance bound by selecting the best model based on validation results from the target dataset [33]. As with prior analyses, OOD performance reflects the best result, regardless of the source domain used for training. We present the results in Fig. 5, where the capacity of the HTR models (measured in millions of parameters) is compared to their OOD performance across the three model selection methods using real and synthetic data.